OpenStudy (anonymous):

Calculus II: 10.8/Maclaurin Series Find Taylor polynomials of orders 0,1,2, and 3 generated by f(x)=(X)^(1/2) at a=4

OpenStudy (anonymous):

Plug into the formula and go out to the thrid derivative when you expand\[[\sum_{0}^{\infty}f \prime ^{n}\left( a \right)\left( x-a \right)^{n}]\div n!\]

OpenStudy (anonymous):

an idea